Find the present value today of $9800 at the end of 8 years using a nominal annual rate of interest of 7% compounded semiannually.
Answer #1

A = P( 1 + i)n

Where

A = 9800

i = 0.07/2

n = 8 * 2 = 16

9800 = P ( 1 + (0.07/2) )16

P = 5651.72